A young girl standing on a bridge throws a stone with an initial velocity of 12 m/s at a downward angle of 45° to the horizontal, in an attempt to hit a block of wood floating in the river below. If the stone is thrown from a height of 20 m and it just reaches the water when the block is 13 m from the bridge, does the stone hit the block?





1
Expert's answer
2022-03-15T10:36:45-0400

Solution.

"v_0=12 m\/s;"

"\theta=45^o;"

"v_{x0}=v_ocos45^o=8.5 m\/s;"

"v_{y0}=-v_osin45^o=-8.5 m\/s;"

"y=-20m;"

"x_b=13m; (x_0=y_0=0);"

"v_y^2=v_{y0}^2-2gh;"

"v_y=\sqrt{144-2\sdot9.81\sdot(-20)}=-22m/s;"

"v_y=v_{0y}-gt\implies t=\dfrac{v_{y0}-v_y}{g};"

"t=\dfrac{-8.5-(-22)}{9.81}=1.4s;"

"x_{max}=v_{0x}t=8.5\sdot1.4=12m;"

Answer: The girl's throw falls short by a meter( the block is at 13m).